You're studying genotypes at a gene that you've discovered contributes to shoulder inflammatory disease. The gene has two alleles: Q and q. QQ homozygotes produce, on average, 2.2 kids. Qq heterozygotes also produce, on average, 2.2 kids. However, qq homozygotes are more likely to have the disease, and therefore, produce, on average, 1.7 kids. What is the relative fitness of the qq genotype? show your work
Answer #1

The genotype with the largest survival rate is known as fittest and it is used as standard for the relative fitness (W) of other groups.

In this question,

QQ and Qq have the genotype frequnency of 2.2 each, therefore are fitter than qq.

In determining the values of relative fitness of all the genotypes (W), all the survival rates are divided by 2.2.

Calculations:

WQQ=2.2/2.2=1

WQq=2.2./2.2=1

Wqq=1.7/2.2=0.77

Therefore, the relative fitnes of qq homozygous kids showing the shoulder inflammatory disease, is 0.77 as compared to 1, in case of QQ homozygous and Qq heterozygous kids.